### v1

# Efficient Preimage Approximation for Neural Network Certification

Abstract
The growing reliance on artificial intelligence in safety- and security-critical applications demands effective neural network certification. A challenging real-world use case is certification against “patch attacks”, where adversarial patches or lighting conditions obscure parts of images, for example traffic signs. One approach to certification, which also gives quantitative coverage estimates, utilizes preimages of neural networks, i.e., the set of inputs that lead to a specified output. However, these preimage approximation methods, including the state-of-the-art PREMAP algorithm, struggle with scalability. This paper presents novel algorithmic improvements to PREMAP involving tighter bounds, adaptive Monte Carlo sampling, and improved branching heuristics. We demonstrate efficiency improvements of at least an order of magnitude on reinforcement learning control benchmarks, and show that our method scales to convolutional neural networks that were previously infeasible. Our results demonstrate the potential of preimage approximation methodology for reliability and robustness certification.
[…]
3853
1 Introduction
The widespread deployment of AI technologies places significant demands on their trustworthiness and reliability, which are of particular concern in high-stakes applications, as recognized in the EU AI Act [9]. Much effort has been devoted to ensuring safety and security of AI applications involving neural networks (NNs) because of their instability to adversarial perturbations [12]. A particularly challenging real-world example are “patch attacks” on images, such as traffic signs. In Figure 1, we show two examples of physical patch attacks: real graffiti on a traffic sign [10] (left) and a patch caused by a sunbeam [5] (middle), which can result from a specific adversarial manipulation of the image or for natural reasons.
Figure 1: Examples of physical patch attacks. Left: real graffiti [10]. Middle: sunbeam [5]. Right: our abstraction of a patch attack.
The above examples highlight the importance of studying robustness of image recognition models to adversarial patch attacks for applications such as safe autonomous driving or biometric security software. To this end, a wide range of methodologies have been developed for neural networks to enhance their robustness, either via (certified) adversarial training [26, 20] or designing certifiable [33] or empirical [31] defences. An alternative method is to certify robustness by computing deterministic [13, 15] or probabilistic [30, 36] guarantees on the network outputs. A drawback of certifiable defences is their relatively high inference time, although there are viable trade-offs [33, 32], Similarly to formal verification, certification is more powerful than adversarial training or empirical defences because it can guarantee that the network output is correct for all inputs in a given region. Using the patch attack in Figure 1 (right) as an illustration, and assuming the patch is placed in a fixed position, certification aims to guarantee the prediction for all possible (RGB) colour values.
Adversarial robustness certification methods commonly work by bounding the outputs of the neural network for a given input subdomain, which may yield loose bounds for the resulting output over-approximation. A recently developed alternative approach focuses instead on bounding the preimage of a given output specification (typically a polyhedron), either through over-approximating the preimage [16, 39] or under-approximating it [38, 39]. Preimage under-approximation, in contrast to over-approximation, offers stronger compliance guarantees, since all inputs in the computed approximation satisfy the guarantee, and quantitative verification, i.e., estimating the proportion of inputs that meet the specification, which can directly serve as a reliability certificate. However, despite the positive experimental results reported in [38, 39], existing methods do not scale to high dimensions, such as commonly used convolutional neural networks, which is compounded by the prohibitive size of the branching computation tree.
In this paper, we aim to significantly improve the scalability and computational performance of the state-of-the-art preimage approximation algorithm PREMAP [38, 39]. This raises a number of challenges, which we resolve by careful analysis that allowed us to select effective algorithmic improvements such as bound tightening, optimized branching and refinement, as well as adaptive Monte Carlo sampling. We experimentally evaluate our method on a range of case studies, showing that our method is able to complete almost twice the number of cases solved using the original method. Now we are able to certify robustness to patch attacks for convolutional networks applied on GTSRB (traffic signs), SVHN (house numbers) and CIFAR (standard image benchmark), which were not feasible with the original PREMAP. We also provide a comparison with the previous version of PREMAP on controller benchmarks, showing at least an order of magnitude improvements in running time. Additionally, we demonstrate how PREMAP can be used for interpretability, to investigate important regions of the input. Our results demonstrate the potential of preimage approximation methodology for exploitation in reliability and robustness certification.
2 Background
This section presents the relevant background. In Section 2.1, we define preliminary concepts and provide an overview of the PREiMage APproximation (PREMAP) method [38, 39], focusing on key aspects needed for our algorithmic improvements; for more detail of the techniques please refer to [39]. Related works are discussed in Section 2.2.
2.1 Preliminaries
Neural networks.
We use $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ to denote a feed-forward neural network, where $f(\mathbf{x})=\mathbf{y}$. For layer $l$, $\mathbf{W}^{l}$ denotes the weight matrix, $\mathbf{b}^{l}$ the bias, $\mathbf{z}^{l}$ the pre-activation neurons, and $h^{l}(\mathbf{z}^{l})$ the activation function, such that $\mathbf{z}^{l}=\mathbf{W}^{l}h^{l-1}(\mathbf{z}^{l-1})+\mathbf{b}^{l}$. We use under- and overlines to denote the lower and upper bounds, e.g., $\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}$. In this paper we focus on ReLU neural networks, where $h^{l}(\mathbf{z}^{l})=\text{ReLU}(\mathbf{z}^{l})=\max(\mathbf{z}^{l},0)$ is applied element-wise, but our method can also be generalized to other activation functions that can be bounded by linear functions [37, 6].
Linear relaxation.
Linear relaxation is used to transform non-convex neuron constraints into linear programs, which provide effective means to approximate a neural network to ensure tractability of analysis. As shown in Figure 2, given concrete lower and upper bounds $\underline{\mathbf{z}}^{l}\leq\mathbf{z}^{l}\leq\overline{\mathbf{z}}^{l}$ on the pre-activation values of layer $l$, there are three cases to consider. When $\overline{\mathbf{z}}^{l}_{i}\leq 0$ then $h^{l}(\mathbf{z}^{l}_{i})=0$ and the neuron is inactive. Similarly, when $\underline{\mathbf{z}}^{l}_{i}\geq 0$ then $h^{l}(\mathbf{z}^{l}_{i})=\mathbf{z}^{l}_{i}$ and the neuron is active. Otherwise, the neuron is unstable and can be bounded by
$\alpha\mathbf{z}^{l}_{i}\leq h^{l}(\mathbf{z}^{l}_{i})\leq(\mathbf{z}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}){\overline{\mathbf{z}}^{l}_{i}}/{(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})},$ (1)
where $\alpha$ is any value in $[0,1]$.
[…]
Figure 2: Linear bounds for inactive, active and unstable ReLU neurons.
For the neural network $f$, linear relaxation is used to compute linear lower and upper bounds for the output like
$\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\leq f(\mathbf{x})\leq\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}},$ (2)
for a given bounded input region. These methods are known as Linear Relaxation based Perturbation Analysis (LiRPA [34]) algorithms.
Preimage approximation.
Given a subset $\mathbb{O}\subset\mathbb{R}^{m}$, the preimage of a function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is defined to be the set of all inputs $\mathbf{x}\in\mathbb{R}^{d}$ that are mapped by $f$ to an element of $\mathbb{O}$. Since computing the exact preimage is intractable, we are interested in computing an under- or over-approximation of the preimage.
To apply LiRPA, we need to bound the input domain $\mathbb{I}\subset\mathbb{R}^{d}$, for example as an axis-aligned hyper-rectangle $\mathbb{I}\subseteq\{\mathbf{x}\in\mathbb{R}^{d}\mid\bigwedge_{i=1}^{d}\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}\}$. We specify the output domain $\mathbb{O}$ as a set of linear inequalities (a polytope) on the output $\bigwedge_{j=1}^{K}(\mathbf{c}_{j}^{\top}\mathbf{y}+d_{j}\geq 0)$. Since the inequalities are linear, they can be implemented by adding a linear layer to the neural network $f_{\mathbb{O}}(\mathbf{x})_{j}=\mathbf{c}_{j}^{\top}f(\mathbf{x})+d_{j}$. Then, by bounding $f_{\mathbb{O}}$ with LiRPA, such that $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\leq f_{\mathbb{O}}(\mathbf{x})\leq\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}$, we obtain an under-approximation $\mathbf{P}$ of the preimage as
$\mathbf{x}\in\mathbb{I}\wedge\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\geq 0\implies f(\mathbf{x})\in\mathbb{O}$
and an over-approximation $\mathbf{P}$ as
$\mathbf{x}\in\mathbb{I}\wedge f(\mathbf{x})\in\mathbb{O}\implies\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}\geq 0.$
PREMAP.
The anytime PREMAP algorithm [39, 38], implements both under- or over-approximation of the preimage of a given output specification $\mathbb{O}$ restricted to the input domain $\mathbb{I}$, in the form of a disjoint union of polytopes. However, in this paper we will focus on under-approximations because they offer stronger guarantees and enable quantitative verification. Algorithm 1 shows an overview of the method with our modification highlighted with green shading (discussed further in Section 3). PREMAP relies on linear relaxation with LiRPA to approximate the preimage (see 3), which is further refined through splitting the problem into disjoint subregions (on 8 and 9), which can be approximated independently in parallel (see 10 and 11) with a divide-and-conquer approach. The algorithm terminates upon reaching a preset threshold for the preimage approximation, with exact rather than statistical quantitative guarantees.
Branching and refinement.
Since the initial bound given by LiRPA might be very loose (Algorithm 1, 3), PREMAP iteratively partitions a subdomain $\mathbb{I}_{\text{sub}}\subseteq\mathbb{I}$ into two smaller subdomains $\mathbb{I}_{\text{sub}}^{-}$ and $\mathbb{I}_{\text{sub}}^{+}$ (5 to 12). Which subdomain to split is selected by the largest difference in estimated volume between the exact preimage and the approximation in the subdomain (lines 6). The split can be either on the input, which partitions the input into two subdomains along some feature, or based on the pre-activation value of an intermediate unstable neuron [39], called ReLU splitting. As this paper deals with larger input domains, for which input splitting performs less well, we will focus on the latter (8).
A key advantage of ReLU splitting is that it allows us to replace unstable neuron bounds with precise bounds. For an unstable ReLU neuron $h^{l}(\mathbf{z}^{l}_{j})=\max(0,\mathbf{z}^{l}_{j})$, we use linear relaxation to bound the post-activation value as in Equation 1. When a split is introduced, the neuron becomes stable in each subdomain, and the exact linear functions $h^{l}(\mathbf{z}^{l}_{j})=\mathbf{z}^{l}_{j}$ and $h^{l}(\mathbf{z}^{l}_{j})=0$ can be used in place of its linear relaxation. This can typically tighten the approximation on each subdomain as the linear relaxation errors for this unstable neuron are removed and substituted with the exact symbolic function.
The bounds, $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}$, for the new under-approximations are optimized via projected gradient descent (Algorithm 2, 3) and the process of refining the subdomains continues until an appropriate termination criterion is met (Algorithm 1, 7). The procedure can also be stopped anytime to produce a valid (but smaller) approximation. The (anytime) union of the preimage approximations $\mathbf{P}$ for the disjoint subdomains $\mathbb{I}_{\text{sub}}$ constitutes the preimage approximation of the output property.
Algorithm 1 Premap is an algorithm for finding preimage under approximations. The inputs are the neural network $f$, the input domain $\mathbb{I}$, the output specifications $\mathbb{O}$, the early stopping threshold $v_{\text{th}}$, and the number of samples $n$. Some functions are detailed in Algorithm 2. The lines with our modifications compared to [39] are highlighted with green shading and are discussed in detail in Sections 3.2, 3.1 and 3.3.
1:function Premap($f,\mathbb{I},\mathbb{O},v_{\text{th}},n$)
2:     $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I},5n)$$\rangle$ ⟩ Sample extra when it is easy
3:     $\mathbf{P}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I},\mathbf{X})$
4:     $\mathbb{D}\leftarrow\{(\mathbf{P},\mathbb{I},\mathbf{X})\}$$\rangle$ ⟩ Approximation, domain, samples
[…]
6:         $\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X}\leftarrow\textsc{Pop}(\mathbb{D},\operatorname*{arg\,max}_{j}(\textsc{Priority}(\mathbb{D}_{j},f_{\mathbb{O}})))$
7:         $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I}_{\text{sub}},n,\mathbf{X})$$\rangle$ ⟩ Top up samples
8:         $l,i\leftarrow\textsc{SelectNeuron}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X})$
9:         $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\textsc{SplitNeuron}(\mathbb{I}_{\text{sub}},\mathbf{X},l,i)$
10:         $\mathbf{P}^{-}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-})$
11:         $\mathbf{P}^{+}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})$
12:         $\mathbb{D}\leftarrow\mathbb{D}\cup\{(\mathbf{P}^{-},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}),\leavevmode\nobreak\ (\mathbf{P}^{+},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})\}$
13:     return $\{\mathbf{P}\mid(\mathbf{P},\cdot,\cdot)\in\mathbb{D}\}$
Algorithm 2 Additional functions for Premap, see Algorithm 1 for details. The LIRPAfunction produces linear bounds using the LiRPA library [34] with optimisable parameters $\boldsymbol{\alpha}$.
1:function PreimageApprox($f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X}$)
2:     $\underline{\mathbf{A}}^{(\boldsymbol{\alpha})},\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}\leftarrow\textsc{LIRPA}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}})$
3:     $\boldsymbol{\alpha},\beta\leftarrow\operatorname*{arg\,max}\limits_{\boldsymbol{\alpha}}\sum\limits_{\mathbf{x}\in\mathbf{X}}\sigma(-\log(\sum\exp({-\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}-\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}})))$
4:     $\mathbf{P}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}+\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}\geq 0\}$
5:     return $\mathbf{P}$$\rangle$ ⟩ Preimage under approximation
6:function EstimateCoverage($\mathbb{D},f_{\mathbb{O}}$)
7:     $v_{\mathbf{P}}\leftarrow\sum_{(\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X})\in\mathbb{D}}\sum_{\mathbf{x}\in\mathbf{X}}\left[\mathbf{x}\in\mathbf{P}\right]/|\mathbf{X}|\cdot|\mathbb{I}_{\text{sub}}|$
8:     $v_{\mathbb{O}}\leftarrow\sum_{(\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X})\in\mathbb{D}}\sum_{\mathbf{x}\in\mathbf{X}}\left[f_{\mathbb{O}}(\mathbf{x})\geq 0\right]/|\mathbf{X}|\cdot|\mathbb{I}_{\text{sub}}|$
9:     return $v_{\mathbf{P}}/v_{\mathbb{O}}$$\rangle$ ⟩ Approximated fraction of the preimage
10:function Priority($\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X},f_{\mathbb{O}}$)
11:     $v_{\mathbb{O}\setminus\mathbf{P}}\leftarrow\sum_{\mathbf{x}\in\mathbf{X}}\left[f_{\mathbb{O}}(\mathbf{x})\geq 0\wedge\mathbf{x}\notin\mathbf{P}\right]/|\mathbf{X}|\cdot\left|\mathbb{I}_{\text{sub}}\right|$
12:     return $v_{\mathbb{O}\setminus\mathbf{P}}$$\rangle$ ⟩ Volume of preimage not approximated
2.2 Related works
Robustness certification methods focus on computing deterministic [13, 15] or probabilistic/statistical [30, 36] guarantees on the output of a neural network. The methods can be classified into complete methods, such as constraint solving [15], or sound though incomplete methods, e.g., convex relaxation [25], which can be strengthened to ensure completeness by employing branch-and-bound procedures [3, 4]. Building on [38, 39], our work employs convex (linear) relaxation and Monte Carlo sampling in a divide-and-conquer framework for approximating the preimage of the neural network with exact (not statistical) quantitative guarantees. We leverage LiRPA [34] as implemented in CROWN tools [37, 35, 29] to compute symbolic relaxation bounds, adapting to preimage approximation. We note that exact preimage computation is intractable for high dimensions, although a variant of this problem known as backward reachability has been studied in control [28] through exact computation or (guaranteed) over-approximation [24, 8]. The method of [16] is limited to preimage over-approximation and scales less well than PREMAP, and consequently our method. Compared to [3, 4], our divide-and-conquer approach focuses on minimizing the difference between the under-approximation and the preimage rather than maximizing a function value.
Methods of defending against patch attacks [31, 32] include active defences, such as small receptive fields or masking out all adversarial pixels from the input image, and passive defences, for example certifiable training using bound propagation. The most common of these, small receptive fields, extract features and predictions from small regions of the image, to avoid spatially localized attacks. The prediction involves robust aggregation such as majority voting. The smaller the receptive fields the better robustness but the worse clean performance, and [33] offers a computationally efficient trade-off.
Compared to active defences, which extend the model, our approach computes guarantees for the underlying neural network models. Certifiable training [19] involves estimating and propagating the bounds of neuron activations in each layer to bound the attacker’s influence on final predictions and is computationally expensive; in contrast, our method approximates the set of inputs for which the prediction is guaranteed for a given trained network.
3 Methods
The aim is to apply PREMAP [38, 39] on larger models with larger input spaces. This means we need to revisit some of the design decisions that worked well on smaller problems. In Section 3.1 we calculate tighter bounds for intermediate layers, in Section 3.2 we describe how we adaptively draw more Monte Carlo samples, and in Section 3.3 we design new heuristics for the selection of which neuron to split next.
3.1 Bound tightening
When a neuron is split, the linear relaxation error from that neuron towards all the following layers is reduced. However, $\beta$-CROWN [29], on which PREMAP is built, typically only updates the final layer because the optimization is computationally intensive [29]. Instead of using the full $\beta$-CROWN optimization, in this work, we propagate the bounds using backwards mode LiRPA [34]. These bounds are not as tight, but much faster to compute.
While LiRPA [34] only propagates bounds to the following layers in the network, a split also implies restrictions in previous layers. Hence, we derive how the bounds of previous layers and the input might tighten after a split. As with backwards mode LiRPA, these bounds are fast to compute but result in looser bounds.
After splitting the $i$:th neuron on layer $l$, we use LiRPA [34] as in Equation 2 to calculate the linear approximation bounds with respect to the preceding layers $m<l$ (including the input) such that
$\underline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\underline{\mathbf{b}}^{lm}\leq\mathbf{z}^{l}\leq\overline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\overline{\mathbf{b}}^{lm}.$
The two constraints defined by the split, $\mathbf{z}^{l}_{i}<0$ and $\mathbf{z}^{l}_{i}\geq 0$, can be propagated backwards using the linear approximations:
$[\underline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\underline{\mathbf{b}}^{lm}]_{i}<0\textrm{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }0\leq[\overline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\overline{\mathbf{b}}^{lm}]_{i}.$ (3)
From this we can update the bounds $\underline{\mathbf{z}}^{m}$ and $\overline{\mathbf{z}}^{m}$ so that:
$\underline{\mathbf{z}}^{m}_{j}\leftarrow\max(\underline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \underline{\mathbf{A}}^{lm}_{ij}<0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow\min(\overline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \underline{\mathbf{A}}^{lm}_{ij}>0$
for $\mathbf{z}^{l}_{i}<0,$ where $c=\sum\nolimits_{k\neq j}\min(\underline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\underline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$ and
$\underline{\mathbf{z}}^{m}_{j}\leftarrow\max(\underline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \overline{\mathbf{A}}^{lm}_{ij}>0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow\min(\overline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \overline{\mathbf{A}}^{lm}_{ij}<0$
for $\mathbf{z}^{l}_{i}\geq 0,$ where $c=\sum\nolimits_{k\neq j}\max(\overline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\overline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$. Note that the division introduces additional numerical instability, which we mitigate with padding around zero, making the bounds slightly looser. After the “reverse” propagation of bounds we use LiRPA to update the bounds of subsequent layers, as mentioned above. This happens in Algorithm 1 at the end of the SplitNeuron function on 9.
3.2 Sampling
The samples are used in PREMAP for three things: Monte Carlo estimation of the volume of both the preimage and the approximation (in Algorithm 2, 7 to 11), the gradient-based optimization of the bounds in $\beta$-CROWN to maximise the preimage approximation volume [39] (on 3 of Algorithm 2), and in the heuristics to select which neuron to split (in Algorithm 1, 8).
We could, like the previous version of PREMAP [39], only draw a large, fixed set of initial samples. Each split defines two subsets where we can reuse the samples (see 9 in Algorithm 1). However, the size of the subsets decrease exponentially with each split. Hence, in this work, when the subsets get too small we draw additional samples (7). At first, we use rejection sampling from uniform samples, $\mathbb{U}(\underline{\mathbf{x}},\overline{\mathbf{x}})$. However, just as with subsets, the hit rate decreases exponentially.
The inequalities introduced in Equation 3 create a polytope in the input space. To sample from this polytope we use hit-and-run sampling [21, 7]. The idea behind hit-and-run sampling is to start from a point inside the polytope, conveniently provided by 9 of Algorithm 1. Then we draw a random direction, calculate the distance to the edge of the polytope, and uniformly sample a new point on this line. Since consecutive steps are correlated, we take multiple steps between every recorded sample. This yields uniformly distributed samples [21] from the polytope and the procedure can be parallelized and GPU accelerated [7]. Note that Equation 3 is an outer bound, so we still need to do rejection sampling on the activations.
3.3 Heuristics
The heuristic for selecting which neuron to split is condensed in Algorithm 1 on 8. What we do is calculate a score for each neuron $i$ on layer $l$ and select the neuron with the highest score. The original PREMAP [39] aims to keep the splits balanced in terms of samples with a score function of
$1-\left|2\sum\nolimits_{i=1}^{n}[\mathbf{z}^{l}_{ij}\geq 0]/n-1\right|,$
where $\mathbf{z}_{ij}^{l}$ is the pre-activation for sample $j$ (out of $n$) at neuron $i$ on layer $l$. However, this heuristic offers no indication of how well LiRPA [34] might tighten the bounds. Therefore, we suggest replacing it with a combination of heuristics described below.
Since the layer $l$ is likely to be in the middle of the network, after a split the negative side, $\mathbf{z}^{l}_{j}<=0.0$, will be constant (by the definition of ReLU) while the positive side will be non-linear with respect to the output. Hence, as heuristics, we estimate how much the split would reduce the “uncertainty” on the negative part.
The heuristics are visualized in Figure 3. First, we calculate the linear bound with respect to the output, $f_{\mathbb{O}}(\mathbf{x})\geq\underline{\mathbf{A}}^{l}\mathbf{z}^{l}+\underline{\mathbf{b}}^{l}$ (or $f_{\mathbb{O}}(\mathbf{x})\leq\overline{\mathbf{A}}^{l}\mathbf{z}^{l}+\overline{\mathbf{b}}^{l}$ for an over-approximation). Then we calculate the area (yellow) that will be reduced,
$\sum_{k}\left|\underline{\mathbf{A}}^{lk}_{i}\underline{\mathbf{z}}^{l}_{i}\underline{\mathbf{z}}^{l}_{i}\right|/\max_{l^{\prime}i^{\prime}}\left(\sum_{k}|\underline{\mathbf{A}}^{l^{\prime}k}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}|\right),$
the maximum under-approximation (green),
$\sum_{k}\left|\underline{\mathbf{A}}^{lk}_{i}\underline{\mathbf{z}}^{l}_{i}\right|/\max_{l^{\prime}i^{\prime}}\left(\sum_{k}|\underline{\mathbf{A}}^{l^{\prime}k}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}|\right),$
and the average extra uncertainty for the samples (blue),
$\sum_{k}\left(\sum_{j=1}^{n}\left|\underline{\mathbf{A}}^{lk}_{i}\min(\mathbf{z}^{l}_{ij},0)\right|/\sum_{j=1}^{n}\left[\mathbf{z}^{l}_{ij}<0\right]\right)/\max_{l^{\prime}i^{\prime}}(\ldots),$
where $k$ represents the, potentially, multiple outputs of $f_{\mathbb{O}}$. We combine these heuristics with a weighted sum that is empirically determined in Section 4.2. Before settling on these heuristics we also considered other options, which are documented in Appendix A.
$\mathbf{z}^{l}_{i}$$f_{\mathbb{O}}(\mathbf{x})^{k}$$\underline{\mathbf{z}}^{l}_{i}$$\overline{\mathbf{z}}^{l}_{i}$$\underline{\mathbf{A}}^{lk}_{i}\mathbf{z}^{l}_{i}$
Figure 3: Visualising the three new neuron selection heuristics. They are the yellow area, the green distance, and the average blue length.
It is possible that the selected split assigns all samples to one branch. This means either that (a) the other side is impossible to reach with the current constraints but the bounds given by LiRPA are not tight enough to show that, or (b) the other side is reachable, but our current estimate for that volume is zero. Hence, we can avoid processing that branch (in Algorithm 1, 10 or 11) without significantly affecting the volume of the under-approximation.
3.4 Parallelisation
In the original PREMAP [39, 38] generating the under- or over-approximation after a split (see Algorithm 1, 10 and 11) is done in parallel. Furthermore, $\beta$-CROWN [29] processes the branches (the while loop on 5) in batches to take advantage of GPU resources. For the pseudocode we opted not to show either for simplicity, but in practice we use both strategies.
Due to the early stopping criteria introduced by PREMAP (5 of Algorithm 1) batching has an additional advantage. If it is difficult to find tighter bounds for the primary branch selected by the heuristics (6 and 8), the “beam search” introduced by the batching might still progress and stop after fewer splits.
4 Experiments
In this section we empirically evaluate our improvements and demonstrate use-cases for PREMAP. In Section 4.1 we describe the experimental setup and datasets and in Section 4.2 we choose the weights for the new heuristics. We compare to the previous version of PREMAP in Section 4.3 and different network architectures in Section 4.4. In Section 4.5 we perform an ablation study and in Section 4.6 we use PREMAP as an investigation tool to enhance interpretability.
4.1 Experimental setup
The experiments were run with four cores from an Intel Xeon 6252 CPU and one NVIDIA RTX 2080 Ti, except for the control tasks in Section 4.3 that were run on an Apple M4 Pro CPU. All preimage approximations were limited to 10 minutes with a batch size of 2 (except the previous version of PREMAP that only processes the splits in parallel). Unless otherwise mentioned, we use a stopping volume threshold of $0.9$.
For most of the experiments we use the following image datasets: GTSRB [27] with traffic signs, SVHN [22] with house numbers, and CIFAR-10 [17] with general images. Details about the model architectures (fully connected and convolutional) and training is in Appendix B. For experiments with image datasets ($32\times 32$ pixels) we randomly sample images, evenly from every class, and patches with random sizes between $3\times 3$ and $6\times 6$ and random positions within the images. Additionally, in Section 4.3, we use five reinforcement learning control tasks further detailed in [39].
The code for the method and experiments will be released under an open source licence upon the acceptance of this paper.
4.2 Parameters
As described in Section 3.3 we now have three heuristics to combine. All of them are normalized to $[0,1]$ to make a weighted sum easier. For this experiment we sample images and patches as described in Section 4.1 and uniformly sample weights for the heuristics. The weights are sampled and evaluated multiple times so that we can normalize the times (by dividing by the mean). We then fit a Gaussian process with an RBF kernel and a white kernel and evaluate it on a grid. The top results can be seen in Table 1, where the “under” heuristic is slightly less important than “area” and “extra”.
Table 1: Using a Gaussian process to find weights for the heuristics. Showing the top results from a grid of $\{0.0,0.25,0.5,0.75,1.0\}$.
Time (s) area under extra
$0.715\pm 0.399$ $1.000$, $0.750$, $1.000$
$0.715\pm 0.400$ $1.000$, $1.000$, $1.000$
$0.718\pm 0.399$ $1.000$, $0.500$, $1.000$
$0.724\pm 0.399$ $1.000$, $0.250$, $1.000$
$0.726\pm 0.399$ $0.750$, $1.000$, $1.000$
4.3 Comparison
Here we compare our version of PREMAP with the previous version [39]. In Table 2 we consider the five smaller reinforcement learning tasks from [39], but run them with ReLU splitting instead of input splitting. In all but dubinsrejoin, the improved splitting heuristic leads to finishing within time and with fewer splits. The cartpole and dubinsrejoin preimages are harder to approximate without input splitting [39], but our new version makes more progress and processes more splits in the same time.
Table 2: Comparing our (new) version of PREMAP to the previous (old) version on five reinforcement learning controllers, averaging ten different seeds. The time limit is ten minutes with a stopping threshold of $0.9$ for the first two datasets and $0.75$ for the last three. We measure the coverage of the under-approximation (larger is better), the number of subdomains required to reach that (smaller is better), and the time it took (smaller is better).
Coverage Domains Time (s)
Dataset new old new old new old
auto_park $0.918$ $0.916$ $2.0$ $3.0$ $0.05$ $0.39$
vcas $0.931$ $0.905$ $17.0$ $21.1$ $0.16$ $5.57$
lunarlander $0.750$ $0.574$ $813.8$ $2745.8$ $36.60$ $600.15$
cartpole $0.750$ $0.424$ $1850.5$ $1601.7$ $118.05$ $600.14$
dubinsrejoin $0.207$ $0.000$ $4742.7$ $1159.0$ $596.11$ $600.16$
Table 3 contains results from the two versions of PREMAP on the GTSRB dataset. We aggregate the results from ten random images from every class with randomly placed and sized patches and count the number of preimage approximations that finish within the time limit. Some preimage approximations are easy enough to finish without entering the loop (Algorithm 1, 5 to 12), which is also visible in Figure 4. If we filter out those results, our version is able to complete more than twice the number of harder cases, especially for the convolutional neural network.
Table 3: Comparing our (new) version of PREMAP with the previous (old) version on the GTSRB dataset. Since some patch and image combinations reach the threshold of $0.9$ in a single iteration, the second row is filtered to exclude those. We measure the average coverage of the under-approximation (larger is better), the number of instances reaching the threshold within 10 minutes (larger is better), and the average time required (smaller is better).
Dataset: GTSRB Coverage Completed Time (s)
Model Subset # new old new old new old
FC All 500 $0.850$ $0.714$ $408$ $351$ $120.55$ $188.37$
FC Hard 292 $0.718$ $0.318$ $174$ $61$ $224.76$ $441.06$
CNN All 500 $0.751$ $0.501$ $358$ $244$ $186.91$ $313.70$
CNN Hard 342 $0.626$ $0.197$ $185$ $54$ $282.32$ $497.88$
4.4 Networks
In Figure 4 we run PREMAP on the image datasets using both fully connected (FC) and convolutional (CNN) neural networks. We divide the results into two groups based on how close to the edge the patch was. Both GTSRB and SVHN show strong dependence on the location of the patch, which is to be expected since the traffic signs and digits are in the centre of the images. Some SVHN images have distracting digits in the surroundings [22] making the robustness at the edges unsurprising.
Figure 4: Time to calculate a preimage under-approximation with a 10 minute time limit and a $0.9$ threshold. The patches are applied to random images (from every class) with random positions and sizes. We group the results based on if the centre of the patch is within 10 pixels of the centre of the image.
[…]
4.5 Ablation
In this section we perform an ablation study on a) the tightening of intermediate bounds and b) the new split selection heuristics. The result can be seen in Figure 5, where we measure the coverage of the under-approximation over time running on random patches, as before, and average the results. Tightening is better than using neither improvement, the heuristics are even better, and using both provides the best results. Note that the tightening of intermediate bounds introduces an additional computational overhead, which reduces the utility on a time versus coverage graph.
Figure 5: Comparing the average approximation coverage (higher is better) over time. The stopping threshold is $0.9$.
Figure 5 also demonstrates the anytime property of the algorithm: if we stop early we still get an under-approximation, just a smaller one, and when given more time more approximations reach the $0.9$ threshold.
4.6 Interpretability
A popular approach to explainable artificial intelligence (XAI) is to cover part of the image to see if the classification changes [14]. This reasoning is easy to understand, since if the prediction changes that area must have been important. However, this requires careful design to make a “neutral” replacement pattern [11], preferably such that the modified image is not be out-of-distribution [2]. With PREMAP we do not consider any singular replacement pattern, but rather the whole preimage of an area.
In Figure 6 we apply $4\times 4$ patches on a grid covering the image on the left. In the second image from the left we show the estimate of the volume of the preimage (not the approximation) for every patch, which indicates that this model is robust against patches of this size on this image. However, in the third image we see that patches around the ‘5’ require more splits for the approximation. This means that in these areas the model is highly non-linear. Furthermore, in the rightmost image we see that the preimage approximation of some patches do not reach a high coverage in the allotted time, notably around the areas where the ‘5’ could be turned into an ‘8’ with two patches. This is how PREMAP can be used to investigate models, which can aid model development and debugging [1].
Figure 6: Using PREMAP to investigate which parts of the image are important for classification without the need to provide a user-defined replacement pattern by applying patches on a grid covering the image. The figures are, from left to right, the original image, the estimated preimage volume (higher means more robust), the number of subdomains required (lower means easier), and the coverage of the preimage under approximations (higher is better).
5 Conclusions
This work introduced algorithmic improvements to the PREMAP algorithm for finding preimage approximations for neural networks. The main enhancements come from the improved heuristics for selecting splits, the calculation of tighter bounds after a split, and the adaptive Monte Carlo sampling. The improved performance is also empirically validated on both smaller control benchmarks and patch attack certification on images. We also demonstrate how to control the trade-off between time and approximation coverage and how PREMAP could be used for explainable AI. While we have focused our experiments on preimage under-approximation because of its inherent advantages for robustness and reliability certification, the methods also extend to over-approximations.
Future work includes applying PREMAP to other domains where certification is needed, for example biometric security or medical diagnosis. PREMAP could also be further enhanced to handle even larger and more complex networks, such as transformers. Finally, extending PREMAP to non-uniform input spaces would unlock additional applications, for example being able to prioritize dangerous conditions.
{ack}
This project received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (FUN2MODEL, grant agreement No. 834115) and ELSA: European Lighthouse on Secure and Safe AI project (grant agreement No. 101070617 under UK guarantee).
Appendix A Heuristics
To improve the heuristic for selecting which neuron to split we consider 12 alternative formulas for calculating the priority. These are detailed in Table 4. Note that not all are new, e.g., "balance" is the one used in the previous version of PREMAP [39, 38] and "lower" [3] and “gap” [23] are used in $\beta$-CROWN [29].
Table 4: Potential heuristics for selecting which neuron $i$ (on layer $l$) to split. There are $j\in[1,\ldots,n]$ samples and $k\in[1,\ldots,K]$ outputs. All formulas not already limited to $[0,1]$ are normalised by dividing by the maximum value and stable neurons with bounds $\underline{\mathbf{z}}^{l}_{i}\geq 0\vee\overline{\mathbf{z}}^{l}_{j}\leq 0$ are skipped.
Heuristic Formula
balance $1-|2\sum_{j}[\mathbf{z}^{l}_{ij}\geq 0]/n-1|$
soft $1-|2\sum_{j}\sigma(\mathbf{z}^{l}_{ij})/n-1|$
lower $\max(-\underline{\mathbf{z}}^{l}_{i},0)$
width $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}$
loose $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}-\bigl{(}\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij})\bigr{)}$
bound $1-\bigl{(}\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij})\bigr{)}/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
gap $(-\underline{\mathbf{z}}^{l}_{i}\cdot\overline{\mathbf{z}}^{l}_{i})/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
area $\sum_{k}|\underline{\mathbf{A}}^{lk}\underline{\mathbf{z}}^{l}_{i}\underline{\mathbf{z}}^{l}_{i}|$
under $\sum_{k}|\underline{\mathbf{A}}^{lk}\underline{\mathbf{z}}^{l}_{i}|$
extra $\sum_{k}\sum_{j}|\underline{\mathbf{A}}^{lk}\min(\mathbf{z}^{l}_{ij},0)|/\sum_{j}[\mathbf{z}^{l}_{ij}<0]$
stable $\min_{j}(\mathbf{z}^{l}_{ij})\geq 0\vee\max_{j}(\mathbf{z}^{l}_{ij})\leq 0$
pure $2\cdot\max\Bigl{(}\left|{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}\geq 0\wedge f(\mathbf{x}_{j})\in\mathbb{O}\bigr{]}}/{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}\geq 0\bigr{]}}-\frac{1}{2}\right|,$
$\left|{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}<0\wedge f(\mathbf{x}_{j})\in\mathbb{O}\bigr{]}}/{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}<0\bigr{]}}-\frac{1}{2}\right|\Bigr{)}$
Some heuristics in Table 4 use the (linear) bounds produced by LiRPA [34, 29]. To minimize computational overhead we store them for reuse after each optimization. Some heuristics use the activation values of the samples. This requires computation on the scale of a forward pass through the network, which is not significant for the overall runtime.
To evaluate the heuristics we draw random samples (i.e., random images with random patches as in Section 4) and run PREMAP multiple times with random combinations of the heuristics and normalize the results by dividing by the mean times for that sample. The weights for the different heuristics are drawn such that roughly half are zero and the rest are uniform.
Figure 7: The effect off using various heuristics for the selection of which neuron to split. The black line is a fitted, single variable, linear regression. Lower values (y) with a non-zero weight (x) indicates that the heuristic helped reduce the number of splits required.
The results for individual heuristics can be seen in Figure 7, where we fit a linear model to more clearly show the trend. Three of the proposed heuristics have a promising slope (negative), with “balance” and “lower” being among the worst (enabling these correlates with longer times). However, some of these heuristics might be correlated, so we cannot choose based on individual plots.
Table 5: Slopes and coefficients of linear models fitted to subsets of the heuristics similar to Figure 7. Lower is better.
Slope Coefficients
$-0.224$ area: $-0.224$
$-0.215$ extra: $-0.215$
$-0.193$ under: $-0.193$
$-0.091$ gap: $-0.091$
$-0.308$ area: $-0.164$, extra: $-0.144$
$-0.303$ area: $-0.179$, under: $-0.124$
$-0.288$ under: $-0.124$, extra: $-0.164$
$-0.224$ gap: $0.000$, area: $-0.225$
$-0.348$ area: $-0.143$, under: $-0.088$, extra: $-0.116$
$-0.287$ gap: $0.045$, area: $-0.175$, extra: $-0.156$
$-0.286$ gap: $0.037$, area: $-0.189$, under: $-0.133$
$-0.274$ gap: $0.031$, under: $-0.132$, extra: $-0.173$
$-0.323$ gap: $0.065$, area: $-0.157$, under: $-0.101$, extra: $-0.131$
$-0.303$ loose: $0.134$, area: $-0.172$, under: $-0.121$, extra: $-0.144$
$-0.302$ width: $0.130$, area: $-0.176$, under: $-0.119$, extra: $-0.137$
$-0.286$ soft: $0.199$, area: $-0.191$, under: $-0.132$, extra: $-0.163$
In Table 5 we enumerate all combinations of the heuristics (only showing the best) and calculate the corresponding linear models. The combination of “area”, “under”, and “extra” continues to impress, since anything else added would have a positive coefficient. This is the selection we use for the paper, further discussed in Sections 3.3 and 4.2.
Appendix B Neural networks
For the image datasets we use two types of neural networks: fully connected and convolutional. The architectures are described in Table 6. The networks were trained with cross-entropy loss using the AdamW [18] optimizer with brightness and hue variation until the accuracy on the test set stopped decreasing (usually around 50 epochs). For the reinforcement learning tasks we used the same networks as in [39].
Table 6: Architectures for the neural networks on the image datasets.
Fully connected Convolutional
Layer Size Layer Size
Input $32\times 32\times 3$ Input $32\times 32\times 3$
Flatten $3072$ Convolutional $30\times 30\times 3$
Linear $300$ ReLU $30\times 30\times 32$
ReLU $300$ Average pooling $15\times 15\times 32$
Linear $300$ Convolutional $12\times 12\times 32$
ReLU $300$ ReLU $12\times 12\times 32$
Linear $300$ Average pooling $6\times 6\times 32$
ReLU $300$ Convolutional $4\times 4\times 64$
Dropout $300$ Average pooling $2\times 2\times 64$
Linear $|\mathbf{y}|$ ReLU $2\times 2\times 64$
Flatten $256$
Dropout $256$
Linear $|\mathbf{y}|$