### v2

# Efficient Preimage Approximation for Neural Network Certification

Abstract
The growing reliance on artificial intelligence in safety- and security-critical applications demands effective neural network certification. A challenging real-world use case is “patch attacks”, where adversarial patches or lighting conditions obscure parts of images, for example, traffic signs. A significant step towards certification against patch attacks was recently achieved using PREMAP, which uses under- and over-approximations of the preimage, the set of inputs that lead to a specified output, for the certification. While the PREMAP approach is versatile, it is currently limited to fully-connected neural networks of moderate dimensionality. In order to tackle broader real-world use cases, we present novel algorithmic extensions to PREMAP involving tighter bounds, adaptive Monte Carlo sampling, and improved branching heuristics. Firstly, we demonstrate that these efficiency improvements significantly outperform the original PREMAP and enable scaling to convolutional neural networks that were previously intractable. Secondly, we showcase the potential of preimage approximation methodology for analysing and certifying reliability and robustness on a range of use cases from computer vision and control. Our software is available in open source.
1 Introduction
The widespread deployment of AI technologies places significant demands on their trustworthiness and reliability, which are of particular concern in high-stakes applications, as recognized in the EU AI Act [1]. Much effort has been devoted to ensuring the safety and security of AI applications involving neural networks (NNs) due to their instability to adversarial perturbations [2]. A particularly challenging real-world example are “patch attacks” on images, such as traffic signs. In Figure 1, we show two examples of physical patch attacks: real graffiti on a traffic sign [3] (left) and a patch caused by a sunbeam [4] (middle).
Figure 1: Examples of physical patch attacks. Left: real graffiti [3]. Middle: sunbeam [4]. Right: our abstraction of a patch attack.
The examples above highlight the importance of studying the robustness of image recognition models to adversarial patch attacks for applications such as safe autonomous driving or biometric security software. To this end, a wide range of methodologies have been developed for neural networks to enhance their robustness, either via (certified) adversarial training [5, 6] or designing certifiable [7] or empirical [8] defences. An alternative method is to certify robustness by computing deterministic [9, 10] or probabilistic [11, 12, 13] guarantees on the network outputs.
Similarly to formal verification, certification is more powerful than adversarial training or empirical defences because it can guarantee that the network output is correct for all inputs in a given region. Using the patch attack in Figure 1 (right) as an illustration, certification aims to guarantee the prediction for all possible (RGB) colour combinations of the pixels in the black square.
Many methods for adversarial robustness certification work by bounding the worst-case output of the neural network for a given input domain. An alternative approach, which gives more information than just the worst-case outputs, focuses on bounding the preimage of a given output specification (typically a polyhedron), either through over-approximating the preimage [14, 15, 16, 17] or under-approximating it [15, 16, 17]. Compared to over-approximation, under-approximation offers stronger compliance guarantees, since all inputs in the computed approximation satisfy the guarantee. Under-approximations also provide quantitative verification, i.e., estimating the proportion of inputs that meet the specification.
Several related methods [14, 17] rely on dividing the input space to obtain tighter approximations, which makes higher-dimensional inputs, such as images, intractable. Furthermore, non-linear activations are the cause of looser bounds [18] and convolutional neural networks (CNN) have significantly more non-linear activations compared to fully connected (FC) networks: even a small image has tens of thousands of non-linear activations in the first layer alone.
In this paper, we aim to broaden the applicability of the state-of-the-art preimage approximation algorithm PREMAP [15, 16] to a wider range of use cases. To this end, we improve the computational performance of PREMAP by significantly reducing the size of the computation tree through calculating tighter bounds, carefully selecting which neuron to branch on, and avoiding heavier optimizations when possible. These improvements enable preimage approximations of convolutional neural networks, which could not be handled previously. We demonstrate through empirical comparison that our method consistently outperforms the original PREMAP [16] on both under- and over-approximations, achieving running time improvements of at least an order of magnitude on reinforcement learning control tasks, while on patch attack certification the time-limited completion rates increase threefold for non-trivial cases.
We also showcase our method on a range of case studies: (i) certification for patch attacks on convolutional networks, which could not be handled previously, applied on GTSRB [19] (traffic signs), SVHN [20] (house numbers) and CIFAR [21] (standard image benchmark); (ii) out-of-distribution (OOD) input detection, where we show how to utilize both under- and over-approximations to certify whether inputs are outliers; and (iii) explainability (XAI), where we demonstrate how PREMAP can complement existing explainability techniques in analysing and debugging neural network models by highlighting important regions in the input. Our results demonstrate the potential of preimage approximation methodology for exploitation in reliability and robustness analysis and certification.
2 Related works
Robustness certification methods focus on computing deterministic [9, 10] or probabilistic/statistical [11, 12] guarantees on the output of a neural network. We can classify the methods into complete methods, such as constraint solving [10], or sound but incomplete methods, e.g., convex relaxation [22], which can be strengthened to completeness by employing branch-and-bound procedures [23, 24].
Building on [15, 16], our work employs convex (linear) relaxation and Monte Carlo sampling in a branch-and-refine framework for approximating the preimage of the neural network with exact (not statistical) quantitative guarantees. We leverage LiRPA [18] as implemented in CROWN tools [25, 26, 27] to compute symbolic relaxation bounds, adapting to preimage approximation.
We note that exact preimage computation is intractable for high dimensions, although a variant of this problem known as backward reachability has been studied in control [28] through exact computation or (guaranteed) over-approximation [29, 30]. Compared to [23, 24], our branch-and-refine approach focuses on minimizing the volume difference between the approximation and the preimage rather than maximizing a function value. The methods of [14, 17] also perform preimage approximations, but the former only over-approximations. However, they both [14, 17] use input-space splitting, which scales poorly beyond low-dimensional data. Methods defending against patch attacks [8, 31] include active defences, such as small receptive fields or masking out all adversarial pixels from the input image, and passive defences, for example, certifiable training using bound propagation. The most common of these, small receptive fields, extract features and predictions from small regions of the image, to avoid spatially localized attacks. The prediction involves robust aggregation such as majority voting. The smaller the receptive fields the better robustness but the worse clean performance, where [7] offers a computationally efficient trade-off. Compared to active defences, which extend the model, our approach computes guarantees for existing neural network models.
Certifiable training [6] involves estimating and propagating the bounds of activations in each layer to bound the attacker’s influence on final predictions and is computationally expensive. Randomized smoothing [13] consists of repeated predictions with random noise added to the input during inference or training [32], which yields probabilistic certification of robustness. In contrast, our method provides quantitative certification of robustness by approximating the set of inputs for which the prediction is guaranteed for a given network.
3 Preliminaries
In this section, we define preliminary concepts and provide an overview of the PREiMage APproximation (PREMAP) method [15, 16], focusing on key aspects needed for our algorithmic improvements; for more detail please refer to [16]
3.1 Neural networks
We use $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ to denote a feed-forward neural network, where $f(\mathbf{x})=\mathbf{y}$. For layer $l$, $\mathbf{W}^{l}$ denotes the weight matrix, $\mathbf{b}^{l}$ the bias, $\mathbf{z}^{l}$ the pre-activation neurons, and $h^{l}(\mathbf{z}^{l})$ the activation function, such that $\mathbf{z}^{l}=\mathbf{W}^{l}h^{l-1}(\mathbf{z}^{l-1})+\mathbf{b}^{l}$. We use under- and over-lines to denote the lower and upper bounds, e.g., $\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}$. In this paper we focus on ReLU neural networks, where $h^{l}(\mathbf{z}^{l})=\text{ReLU}(\mathbf{z}^{l})=\max(\mathbf{z}^{l},0)$ is applied element-wise, but our method can also be generalized to other activation functions that can be bounded by linear functions [25, 33].
3.2 Linear relaxation
Linear relaxation is used to transform non-convex neuron constraints into linear programs, which provide effective means to approximate a neural network to ensure tractability of analysis. As shown in Figure 2, given concrete lower and upper bounds $\underline{\mathbf{z}}^{l}\leq\mathbf{z}^{l}\leq\overline{\mathbf{z}}^{l}$ on the pre-activation values of layer $l$, there are three cases to consider. When $\overline{\mathbf{z}}^{l}_{i}\leq 0$ then $h^{l}(\mathbf{z}^{l}_{i})=0$ and the neuron is inactive. Similarly, when $\underline{\mathbf{z}}^{l}_{i}\geq 0$ then $h^{l}(\mathbf{z}^{l}_{i})=\mathbf{z}^{l}_{i}$ and the neuron is active. Otherwise, the neuron is unstable and can be bounded by
$\alpha\mathbf{z}^{l}_{i}\leq h^{l}(\mathbf{z}^{l}_{i})\leq(\mathbf{z}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}){\overline{\mathbf{z}}^{l}_{i}}/{(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})},$ (1)
where $\alpha$ is any value in $[0,1]$.
[…]
Figure 2: Linear bounds for inactive, active and unstable ReLU neurons.
For the neural network $f$, linear relaxation is used to compute linear lower and upper bounds for the output:
$\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\leq f(\mathbf{x})\leq\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}},$ (2)
for a given bounded input region. These methods are known as Linear Relaxation based Perturbation Analysis (LiRPA [18]) algorithms.
3.3 Preimage approximation
Given a subset $\mathbb{O}\subset\mathbb{R}^{m}$, the preimage of a function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is defined to be the set of all inputs $\mathbf{x}\in\mathbb{R}^{d}$ that are mapped by $f$ to an element of $\mathbb{O}$. Since computing the exact preimage is intractable, we are interested in computing an under- or over-approximation of the preimage.
To apply LiRPA, we need to bound the input domain $\mathbb{I}\subset\mathbb{R}^{d}$, for example as an axis-aligned hyper-rectangle $\mathbb{I}\subseteq\{\mathbf{x}\in\mathbb{R}^{d}\mid\bigwedge_{i=1}^{d}\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}\}$. We specify the output domain $\mathbb{O}$ as a set of linear inequalities (a polytope) on the output $\bigwedge_{j=1}^{K}(\mathbf{c}_{j}^{\top}\mathbf{y}+d_{j}\geq 0)$. Since the inequalities are linear, they can be implemented by adding a linear layer to the neural network $f_{\mathbb{O}}(\mathbf{x})_{j}=\mathbf{c}_{j}^{\top}f(\mathbf{x})+d_{j}$. Then, by bounding $f_{\mathbb{O}}$ with LiRPA, such that $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\leq f_{\mathbb{O}}(\mathbf{x})\leq\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}$, we obtain an under-approximation $\mathbf{P}$ of the preimage as
$\mathbf{x}\in\mathbb{I}\wedge\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}\geq 0\implies f(\mathbf{x})\in\mathbb{O}$
and an over-approximation $\mathbf{P}$ as
$\mathbf{x}\in\mathbb{I}\wedge f(\mathbf{x})\in\mathbb{O}\implies\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}\geq 0.$
3.4 PREMAP
The PREMAP algorithm [16, 15], implements both under- or over-approximation of the preimage of a given output specification $\mathbb{O}$ restricted to the input domain $\mathbb{I}$, in the form of a disjoint union of polytopes. Algorithm 1 shows an overview of the PREMAP method with our modifications highlighted with green shading (discussed further in Section 4). For brevity, some equations are only shown for under-approximations, but work equivalently for over-approximations. PREMAP relies on linear relaxation with LiRPA to approximate the preimage (see line 3), which is further refined through splitting the problem into disjoint subregions (on 8 and 10), that can be approximated independently in parallel (see 12 and 13) with a branch-and-refine approach. The algorithm can be terminated anytime, but stops upon reaching a preset threshold for the preimage approximation.
3.5 Branching and refinement
Since the initial bound given by LiRPA might be very loose (Algorithm 1, line 3), PREMAP iteratively partitions a subdomain $\mathbb{I}_{\text{sub}}\subseteq\mathbb{I}$ into two smaller subdomains $\mathbb{I}_{\text{sub}}^{-}$ and $\mathbb{I}_{\text{sub}}^{+}$ (5 to 14). We select which subdomain to split by the largest difference in estimated volume between the exact preimage and the approximation in the subdomain (line 6). The split can be either on the input, which partitions the input into two subdomains along some feature, or based on the pre-activation value of an intermediate unstable neuron [16], called ReLU splitting. As this paper deals with larger input domains, for which input splitting performs less well, we will focus on the latter (line 8).
A key advantage of ReLU splitting is that it allows us to replace unstable neuron bounds with precise bounds. For an unstable ReLU neuron $h^{l}(\mathbf{z}^{l}_{j})=\max(0,\mathbf{z}^{l}_{j})$, we use linear relaxation to bound the post-activation value as in Equation 1. When a split is introduced, the neuron becomes stable in each subdomain, and the exact linear functions $h^{l}(\mathbf{z}^{l}_{j})=\mathbf{z}^{l}_{j}$ and $h^{l}(\mathbf{z}^{l}_{j})=0$ can be used in place of its linear relaxation. This typically tightens the approximation on each subdomain as the linear relaxation errors for this unstable neuron are removed and substituted with the exact symbolic function.
The bounds, $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}$, for the new under-approximations (or $\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}$ for over-approximations) are optimized via projected gradient descent to maximize the approximation (Algorithm 2, line 3) and the process of refining the subdomains continues until an appropriate termination criterion is met (Algorithm 1, line 7). The procedure can also be stopped anytime to produce a valid (but looser) approximation. The union of the preimage approximations $\mathbf{P}$ for the disjoint subdomains $\mathbb{I}_{\text{sub}}$ constitutes the preimage approximation of the output property.
1:function Premap($f,\mathbb{I},\mathbb{O},v_{\text{th}},n$)
2:  $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I},5n)$ $\rangle\rangle$ Sample extra when it is easy
3:  $\mathbf{P}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I},\mathbf{X})$
4:  $\mathbb{D}\leftarrow\{(\mathbf{P},\mathbb{I},\mathbf{X})\}$ $\rangle\rangle$ Approximation, domain, samples
5:  while $\textsc{EstimateApproxRatio}(\mathbb{D},f_{\mathbb{O}})<v_{\text{th}}$
6:   $\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X}\leftarrow\textsc{Pop}(\mathbb{D},\operatorname*{arg\,max}_{j}(\textsc{Priority}(\mathbb{D}_{j},f_{\mathbb{O}})))$
7:   $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I}_{\text{sub}},n,\mathbf{X})$ $\rangle\rangle$ See Section 4.2
8:   $l,i\leftarrow\textsc{SelectNeuron}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X})$ $\rangle\rangle$ Section 4.3
9:   $\mathbb{I}_{\text{sub}}\leftarrow\textsc{CheckShortcut}(\mathbb{I}_{\text{sub}},\mathbf{X})$ $\rangle\rangle$ See Section 4.4
10:   $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\textsc{SplitNeuron}(\mathbb{I}_{\text{sub}},\mathbf{X},l,i)$
11:   $\textsc{PropagateBounds}(\mathbb{I}_{\text{sub}}^{-},\mathbb{I}_{\text{sub}}^{+},l,i)$ $\rangle\rangle$ See Section 4.1
12:   $\mathbf{P}^{-}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-})$
13:   $\mathbf{P}^{+}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})$
14:   $\mathbb{D}\leftarrow\mathbb{D}\cup\{(\mathbf{P}^{-},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}),~(\mathbf{P}^{+},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})\}$
15:  return $\{\mathbf{P}\mid(\mathbf{P},\cdot,\cdot)\in\mathbb{D}\}$
Algorithm 1: Premap is an algorithm for finding preimage under approximations. The inputs are the neural network $f$, the input domain $\mathbb{I}$, the output specifications $\mathbb{O}$, the early stopping threshold $v_{\text{th}}$, and the number of samples $n$. Some functions are detailed in Algorithm 2. The lines with improvements compared to [16] are highlighted with green shading and are discussed in Sections 4.2, 4.1 and 4.3.
1:function PreimageApprox($f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X}$)
2:  $\underline{\mathbf{A}}^{(\boldsymbol{\alpha})},\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}\leftarrow\textsc{Lirpa}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}})$
3:  $\boldsymbol{\alpha}\leftarrow\operatorname*{arg\,max}\limits_{\boldsymbol{\alpha}}\sum\limits_{\mathbf{x}\in\mathbf{X}}\sigma(-\log(\sum\exp({-\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}-\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}})))$
4:  $\mathbf{P}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}+\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}\geq 0\}$
5:  return $\mathbf{P}$ $\rangle\rangle$ Preimage under approximation
6:function EstimateApproxRatio($\mathbb{D},f_{\mathbb{O}}$)
7:  $v_{\mathbf{P}}\leftarrow\sum_{(\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X})\in\mathbb{D}}\sum_{\mathbf{x}\in\mathbf{X}}\left[\mathbf{x}\in\mathbf{P}\right]/|\mathbf{X}|\cdot|\mathbb{I}_{\text{sub}}|$
8:  $v_{\mathbb{O}}\leftarrow\sum_{(\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X})\in\mathbb{D}}\sum_{\mathbf{x}\in\mathbf{X}}\left[f_{\mathbb{O}}(\mathbf{x})\geq 0\right]/|\mathbf{X}|\cdot|\mathbb{I}_{\text{sub}}|$
9:  return $v_{\mathbf{P}}/v_{\mathbb{O}}$ $\rangle\rangle$ Approximated fraction of the preimage
10:function Priority($\mathbf{P},\mathbb{I}_{\text{sub}},\mathbf{X},f_{\mathbb{O}}$)
11:  $v_{\mathbb{O}\setminus\mathbf{P}}\leftarrow\sum_{\mathbf{x}\in\mathbf{X}}\left[f_{\mathbb{O}}(\mathbf{x})\geq 0\wedge\mathbf{x}\notin\mathbf{P}\right]/|\mathbf{X}|\cdot\left|\mathbb{I}_{\text{sub}}\right|$
12:  return $v_{\mathbb{O}\setminus\mathbf{P}}$ $\rangle\rangle$ Volume of preimage not approximated
13:function SplitNeuron($\mathbb{I}_{\text{sub}},\mathbf{X},l,i$)
14:  $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\mathbf{z}^{l}_{i}<0\},\{\mathbf{x}\mid\mathbf{x}\in\mathbf{X}\wedge\mathbf{z}^{l}_{i}<0\}$
15:  $\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\mathbf{z}^{l}_{i}\geq 0\},\{\mathbf{x}\mid\mathbf{x}\in\mathbf{X}\wedge\mathbf{z}^{l}_{i}\geq 0\}$
16:  return $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}$ $\rangle\rangle$ Split domain and samples
Algorithm 2: Additional functions for Premap, see Algorithm 1 for details. The Lirpa[18] function produces linear bounds with optimizable parameters $\boldsymbol{\alpha}$. And $\mathbf{z}^{l}_{i}$ is the activation of neuron $i$ on layer $l$.
4 Methods
We aim to broaden the applicability of PREMAP [15, 16] to a range of use cases, which necessitates improvements in scalability and efficiency, including an extension to convolutional networks. Therefore, we will revisit some design decisions that worked well on smaller problems. In Section 4.1 we calculate tighter bounds for intermediate layers, in Section 4.2 we describe how we adaptively draw more Monte Carlo samples, in Section 4.3 we design new heuristics to select which neuron to split next, and in Sections 4.4 and 4.5 we describe computational strategies for increased efficiency.
4.1 Bound tightening
When a neuron is split, the linear relaxation error from that neuron to all the following layers is reduced. However, $\beta$-CROWN [27], on which PREMAP is built, typically only updates the final layer because the optimization is computationally intensive [27]. In this work, instead of using full $\beta$-CROWN optimization, for intermediate layers we propagate the bounds using simpler LiRPA [18]. These bounds are not as tight, but much faster to compute.
While LiRPA [18] only propagates bounds to the following layers in the network, a split also implies restrictions in previous layers. Hence, we derive how the bounds of previous layers and the input might tighten after a split.
After splitting the $i$:th neuron on layer $l$, we use LiRPA [18] as in Equation 2 to calculate the linear approximation bounds with respect to the preceding layers $m<l$ (including the input) such that
$\underline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\underline{\mathbf{b}}^{lm}\leq\mathbf{z}^{l}\leq\overline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\overline{\mathbf{b}}^{lm}.$
The two constraints defined by the split, $\mathbf{z}^{l}_{i}<0$ and $\mathbf{z}^{l}_{i}\geq 0$, can be propagated backwards using the linear approximations:
[…]
From this we can update the bounds $\underline{\mathbf{z}}^{m}$ and $\overline{\mathbf{z}}^{m}$ so that:
$\underline{\mathbf{z}}^{m}_{j}\leftarrow$ $\max(\underline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\underline{\mathbf{A}}^{lm}_{ij}<0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow$ $\min(\overline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\underline{\mathbf{A}}^{lm}_{ij}>0$
for $\mathbf{z}^{l}_{i}<0,$ where $c=\sum\nolimits_{k\neq j}\min(\underline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\underline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$ and
$\underline{\mathbf{z}}^{m}_{j}\leftarrow$ $\max(\underline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\overline{\mathbf{A}}^{lm}_{ij}>0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow$ $\min(\overline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\overline{\mathbf{A}}^{lm}_{ij}<0$
for $\mathbf{z}^{l}_{i}\geq 0,$ where $c=\sum\nolimits_{k\neq j}\max(\overline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\overline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$. Note that the division introduces additional numerical instability, which we mitigate with padding around zero, making the bounds slightly looser. After the “reverse” propagation of bounds we use LiRPA to update the bounds of subsequent layers. This happens in Algorithm 1 on line 11.
4.2 Sampling
The samples are used in PREMAP in three ways: for Monte Carlo estimation of the volume of both the preimage and the approximation (on line Algorithm 2, 7 to 11), the gradient-based optimization of the bounds in $\beta$-CROWN to maximize the preimage approximation volume [16] (on line 3 of Algorithm 2), and in the heuristics to select which neuron to split (in Algorithm 1, line 8).
We note that the original PREMAP [16] only draws a large, fixed set of initial samples. Each split defines two subsets where we can reuse the samples (see line 10 in Algorithm 1). However, the size of the subsets decrease exponentially with each split. Therefore, when the subsets get too small, we draw additional samples (line 7). At first, we use rejection sampling, however, the hit rate also decreases exponentially.
The inequalities introduced in Equation 3 create a polytope in the input space. To sample from this polytope we use hit-and-run sampling [34, 35]. The idea behind hit-and-run sampling is to start from a point inside the polytope, conveniently provided by line 10 of Algorithm 1. Then we draw a random direction, calculate the distance to the edge of the polytope, and uniformly sample a new point on this line. Since consecutive steps are correlated, we take multiple steps between every recorded sample. This yields uniformly distributed samples [34] from the polytope and the procedure is GPU accelerated [35]. Note that Equation 3 is an outer bound, so we still need to do rejection sampling on the activations.
4.3 Heuristics
In this section we discuss how to design and select appropriate heuristics to improve computational performance of preimage approximation, both for under- and over-approximations. The heuristic for selecting which neuron to split is condensed in Algorithm 1 on line 8, where a score is calculated for each neuron and the neuron with the highest score is selected. The original PREMAP [16] aims to keep the splits balanced in terms of samples with a score function that counts the number of samples on each side of the split:
$1-\left|2\sum\nolimits_{i=1}^{n}[\mathbf{z}^{l}_{ij}\geq 0]/n-1\right|,$
where $\mathbf{z}_{ij}^{l}$ is the pre-activation for sample $j$ (out of $n$) at neuron $i$ on layer $l$. However, this heuristic offers no indication of how well LiRPA [18] might tighten the bounds.
To improve the selection of which neuron to split, we consider alternative heuristics to calculate priority, detailed in Table I. We include the balance heuristic from the original PREMAP [16, 15] and the lower [23] and gap [36] heuristics used in $\beta$-CROWN [27]. All heuristics are normalized by dividing by the maximum value, unless already in $[0,1]$.
TABLE I: Heuristics for selecting which neuron to split
Heuristic Formula
balance [16] $1-|2\sum_{j}[\mathbf{z}^{l}_{ij}\geq 0]/n-1|$
soft $1-|2\sum_{j}\sigma(\mathbf{z}^{l}_{ij})/n-1|$
lower [23] $\max(-\underline{\mathbf{z}}^{l}_{i},0)$
width $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}$
loose $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}-(\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij}))$
bound $1-(\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij}))/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
gap [36] $(-\underline{\mathbf{z}}^{l}_{i}\cdot\overline{\mathbf{z}}^{l}_{i})/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
area $\sum_{k}|\underline{\mathbf{A}}^{lk}\underline{\mathbf{z}}^{l}_{i}\underline{\mathbf{z}}^{l}_{i}|$
under $\sum_{k}|\underline{\mathbf{A}}^{lk}\underline{\mathbf{z}}^{l}_{i}|$
extra $\sum_{k}\sum_{j}|\underline{\mathbf{A}}^{lk}\min(\mathbf{z}^{l}_{ij},0)|/\sum_{j}[\mathbf{z}^{l}_{ij}<0]$
All formulas are for a single neuron (layer $l$, index ii), with $j\in[1,\ldots,n]$ samples in $\mathbf{X}$ and $k\in[1,\ldots,K]$ outputs of $f_{\mathbb{O}}$.
In Table I the balance [16] and soft heuristics prioritize neurons that would evenly split the samples. The lower [23] and width heuristics measure the bounds given by LiRPA, while loose and bound compare these bounds with the actual minimum and maximum values achieved by the samples. The gap [36] heuristic measures the vertical distance between the linear upper bound and origin in Figure 2.
As visualized in Figure 3, after a ReLU activation negative values $\mathbf{z}^{l}_{j}<=0.0$ are truncated while the positive values are non-linear with respect to the output of $f_{\mathbb{O}}$. Hence, the area, under, and extra heuristics estimate how much a split would reduce the “output uncertainty” on the negative part.
$\mathbf{z}^{l}_{i}$$f_{\mathbb{O}}(\mathbf{z}^{l}_{i})^{k}$$\underline{\mathbf{z}}^{l}_{i}$$\overline{\mathbf{z}}^{l}_{i}$$\underline{\mathbf{A}}^{lk}_{i}\mathbf{z}^{l}_{i}$
Figure 3: Visualizing three neuron selection heuristics: area (yellow area), under (green distance), and extra (the average blue length at the samples).
Some heuristics in Table I use the (linear) bounds produced by LiRPA [18, 27]. To minimize computational overhead we cache them after each optimization. Other heuristics use the activation values of the samples, which requires a forward pass through the network. Hence, no heuristic significantly extends the overall runtime. The selection of which heuristics to use and their weights is performed in Section 5.2.
4.4 Shortcuts
In some cases, a branch can be discarded or finalized before running the heavy $\beta$-CROWN process, for example, when the branch contains no preimage samples (when under-approximating) or only preimage samples (when over-approximating). This lets us avoid some heavy computation (in Algorithm 1, on line 12 or 13) on that branch. For the remaining branch we can perform an additional split before running $\beta$-CROWN.
Furthermore, if all samples are assigned to one branch of a split it means that either (a) the other branch is impossible to reach, but the current LiRPA bounds are not tight enough, or (b) the other side is reachable, but our current estimate for that volume is near zero. We can detect these splits with just a single forward pass through the network, which lets us split and stabilize all of them simultaneously on line 9 of Algorithm 1 (before invoking $\beta$-CROWN).
4.5 Batching
The original PREMAP [16, 15] generates the approximations of a split in parallel (see Algorithm 1, 12 and 13). In addition, $\beta$-CROWN[27] processes the branches (the loop on line 5) in batches to take advantage of GPU resources. The pseudocode shows neither for readability, but in practice we use both strategies.
Due to the early stopping criteria introduced by PREMAP [16, 15] (line 5 of Algorithm 1), batching has an additional advantage: If it is difficult to find tighter bounds for the primary branch selected by the heuristics (6 and 8), the secondary branches in the batch might still progress and reach the threshold earlier.
5 Empirical evaluation
In this section we empirically evaluate the quality of our improvements to PREMAP. In Section 5.1 we describe the experimental setup and datasets and in Section 5.2 we choose the weights for the new heuristics. We compare to the original PREMAP on patch attacks and reinforcement learning tasks in Section 5.3 and on different network architectures in Section 5.4. In Section 5.5 we perform an ablation study.
5.1 Experimental setup
We ran all our experiments (including the use cases in Section 6) with four cores from an Intel Xeon 6252 CPU and one NVIDIA RTX 2080 Ti. All preimage approximations were limited to 10 minutes with a batch size of 2 (except the original PREMAP that does not support batching). Unless otherwise mentioned, we perform under-approximations with a stopping threshold for the approximation ratio of $0.9$ ($1.1$ for over-approximations).
We certify robustness against patch attacks on the following image datasets: GTSRB [19] (traffic signs), SVHN [20] (house numbers), and CIFAR-10 [21] (general images). Details about the model architectures (a fully connected network with three $300$ wide layers and a convolutional network with three convolutional layers and one fully connected layer) and training are included in Appendix B. For experiments with image datasets ($32\times 32$ pixels) we randomly sample images, evenly from every class. The patches are rectangular, with sizes between $3\times 3$ and $6\times 6$, and are randomly placed within the images.
In Section 5.3, we use the same five [37, 38, 39] (low-dimensional) reinforcement learning control tasks and networks from [16]. We compute preimage approximations for select actions (classes), like fire main engine in lunarlander and push left in cartpole, for the same input domains as [16] (with $x_{v}\in[-2,0]$} for lunarlander, $x_{\omega}\in[-2,0]$} for cartpole, and $x_{v}\in[-0.1,0.1]$} for dubinsrejoin).
For the use cases in Section 6 we, additionally, use the same synthetic dataset and network as in [14] for outlier detection and create a new dataset from GTSRB [19] and CIFAR-10 [21] for evaluating explanations.
The code for both the method and the experiments is available under an open source licence [^1] Available at: https://github.com/Anton-Bjorklund/Premap2.
5.2 Hyperparameters
To select which heuristics to use we sample images and patches as described in Section 5.1 and run the reinforcement learning control tasks with different seeds. For each task we uniformly sample multiple weights for the heuristics so that half are inactive and at least one has a weight of one:
$\boldsymbol{w}^{\prime}$ $\sim\mathbb{U}[-1,1]$
$\boldsymbol{w}$ $=\max(\boldsymbol{0},\boldsymbol{w}^{\prime})/\max(\boldsymbol{w}^{\prime}),$
and perform both over- and under-approximations.
To evaluate the heuristics we calculate the difference in the approximation ratio from after the first iteration (Algorithm 1, line 3) to after the final iteration (or 5 minute time limit) and divide by the total time:
$\delta=\left|\frac{v_{\mathbf{P}}(1)}{v_{\mathbb{O}}}-\frac{v_{\mathbf{P}}(n)}{v_{\mathbb{O}}}\right|/t(n).$ (4)
Larger values of $\delta$ indicate that the approximation ratio improves more quickly. To compare across tasks and datasets we normalize the $\delta$ values by dividing by the maximum $\delta$ for each task. We also filter out tasks where the heuristic choice makes no difference, such as when PREMAP finishes after the first iteration.
The results for individual heuristics can be seen in Figure 4, with a linear model showing the general trend. Enabling the “area”, “under”, and “extra” heuristics are correlated with faster improvements to the approximation ratio, while using “balance” [16] and “lower” [23] instead of another heuristic leads to slower approximations.
Figure 4: The effect of using various heuristics for the selection of which neuron to split (each dot represents a different combination). The black line is a single variable, linear regression, where a positive slope is correlated with faster improvements, see Equation 4.
However, some heuristics are correlated and some might complement each other, so we cannot choose the combination based on individual plots. To evaluate how heuristics combine we fit a kernel ridge regression model, using a grid search with cross-validation to set the parameters ($\alpha=8$, $\gamma=1.0$, and a Laplacian kernel). In Table II we evaluate the model on a grid, $\boldsymbol{w}\in\{0.0,0.25,0.5,0.75,1.0\}^{10}$, and show the best results for increasing numbers of non-zero coefficients for the different heuristics.
TABLE II: Comparing different combinations of heuristics
$\boldsymbol{\delta}$ Weights
$0.722$ under: $1.00$
$0.712$ extra: $1.00$
$0.680$ area: $1.00$
$0.510$ bound: $1.00$
$0.810$ extra: $1.00$, area: $0.75$
$0.790$ under: $1.00$, gap: $0.75$
$0.788$ extra: $1.00$, area: $1.00$
$0.786$ extra: $1.00$, area: $0.50$
$0.838$ extra: $1.00$, area: $0.75$, gap: $0.25$
$0.833$ extra: $1.00$, area: $0.75$, gap: $0.50$
$0.823$ extra: $1.00$, area: $0.75$, under: $0.50$
$0.823$ extra: $1.00$, area: $0.75$, under: $0.25$
$0.843$ extra: $1.00$, area: $0.75$, under: $0.50$, gap: $0.25$
$0.842$ extra: $1.00$, area: $0.75$, under: $0.25$, gap: $0.25$
$0.842$ soft: $1.00$, extra: $0.75$, area: $0.75$, under: $0.50$
$0.839$ extra: $1.00$, under: $0.75$, area: $0.75$, gap: $0.50$
$0.839$ soft: $1.00$, extra: $0.75$, area: $0.75$, under: $0.50$, gap: $0.25$
$0.837$ soft: $1.00$, extra: $0.75$, area: $0.75$, under: $0.50$, width: $0.25$
$0.836$ extra: $1.00$, area: $0.75$, under: $0.25$, gap: $0.25$, soft: $0.25$
$0.836$ soft: $1.00$, extra: $0.75$, under: $0.75$, area: $0.75$, gap: $0.25$
For the rest of the experiments in this paper we use the following heuristics and weights: “extra” at $1.0$, “area” at $0.75$, “under” at $0.5$, and “gap” at $0.25$. Separate results for under- and over-approximations can be seen in Appendix A.
5.3 Comparison with original PREMAP
In this section we compare our version of PREMAP with the original PREMAP [16]. In Table III we consider the five reinforcement learning tasks from [16], but run them with ReLU splitting instead of input splitting (note that cartpole and dubinsrejoin are better suited for input splitting [16]). For these tasks we run PREMAP only on the CPU, since the GPU would introduce a lot of overhead due to the small size of the models. Each task is run ten times with different seeds for both under- and over-approximations.
Over-approximations are easier for both versions, and our version is consistently faster for both under- and over-approximations. Our version only exceeds the 10-minute time limit when under-approximating dubinsrejoin, but still makes more progress (higher approximation ratio).
TABLE III: Comparing our version of premap to the original
Dataset Approx. ratio Iterations Time (s)
threshold new old new old new old
auto_park $0.9$ $\leq$ $1.000$ $0.915$ $3.9$ $10.5$ $0.21$ $0.94$
vcas $0.9$ $\leq$ $0.964$ $0.905$ $7.0$ $31.1$ $0.80$ $3.85$
lunarlander $0.75$ $\leq$ $0.750$ $0.488$ $171.8$ $4930.2$ $29.97$ $600.21$
cartpole $0.75$ $\leq$ $0.776$ $0.266$ $15.7$ $4661.5$ $20.55$ $600.21$
dubinsrejoin $0.75$ $\leq$ $0.465$ $0.000$ $1421.4$ $3725.1$ $595.97$ $600.25$
Under-approximation $\uparrow$       $\downarrow$ Over-approximation
auto_park $1.1$ $\geq$ $1.016$ $1.032$ $3.0$ $14.0$ $0.20$ $2.06$
vcas $1.1$ $\geq$ $1.058$ $1.059$ $1.0$ $1.0$ $0.06$ $0.08$
lunarlander $1.25$ $\geq$ $1.195$ $1.232$ $10.5$ $53.4$ $1.55$ $8.40$
cartpole $1.25$ $\geq$ $1.235$ $1.236$ $10.2$ $98.3$ $3.08$ $11.47$
dubinsrejoin $1.25$ $\geq$ $1.238$ $1.250$ $10.1$ $1865.6$ $2.12$ $221.08$
Table IV contains results from the two versions of PREMAP on the GTSRB dataset. This comparison required applying some of the changes to the original PREMAP, so that it can technically run CNNs. Since LiRPA [18] and $\beta$-CROWN[27] already handle convolutions, we have adopted the same strategy for PREMAP. We also made PREMAP agnostic to the additional dimensions in the activation tensors.
TABLE IV: Comparing our version of PREMAP to the original
Dataset: GTSRB Approx. ratio Completed Time (s)
Model Subset new old new old new old
FC All $0.885$ $0.529$ $0.840$ $0.481$ $117.50$ $321.24$
FC Hard $0.811$ $0.287$ $0.757$ $0.215$ $176.83$ $484.21$
CNN All $0.692$ $0.353$ $0.692$ $0.343$ $211.31$ $400.69$
CNN Hard $0.620$ $0.201$ $0.620$ $0.190$ $259.93$ $490.98$
Under-approximation $\uparrow$       $\downarrow$ Over-approximation
FC All $1.014$ $1.079$ $0.920$ $0.810$ $51.00$ $123.86$
FC Hard $1.352$ $1.807$ $0.273$ $0.095$ $446.74$ $573.37$
CNN All $1.011$ $1.107$ $0.983$ $0.829$ $19.12$ $114.12$
CNN Hard $2.486$ $6.764$ $0.818$ $0.000$ $173.38$ $600.00$
Some preimage approximations are easy enough to finish without entering the loop (Algorithm 1, lines 5 to 14), which is also visible in Figure 5. In Table IV we show the results for all cases (rows labelled “All”) and with the trivial results filtered out (rows labelled “Hard”). Our version is able to complete more than thrice the number of harder cases within the 10-minute time limit.
5.4 Comparison on different network architectures
In Figure 5 we show plots for PREMAP on the image datasets using both fully connected and convolutional neural networks. We divide the results into two groups, based on whether the patch is close to the edge of the image.
For under-approximations, both GTSRB and SVHN show strong dependence on the location of the patch, which is to be expected since the traffic signs and digits are in the centre of the images. Some SVHN images even have distracting digits in the surroundings [20], which would further emphasize the centre during training.
The LiRPA processing time and the maximum number of splits is dependent on the number of unstable neurons. For CNNs, the number of activations scales with the size of the image. Here the first layer has $30\times 30\times 32=28~800$ activations! As can be seen in Figure 5, this makes the CNNs slower and harder than the fully connected networks, even though they have fewer parameters.
Figure 5: Time to calculate a preimage under-approximation with a 10-minute time limit and a $0.9$ threshold. The patches are applied to random images (from every class) with random positions and sizes. We group the results based on whether the centre of the patch is within 10 pixels of the centre.
Since most non-adversarial perturbations do not change the predicted class, the preimages tend to be quite large (close to 100% coverage). This makes many over-approximations trivial as the initial domain already has approximation ratio below $1.1$, especially for the more robust CNNs.
5.5 Ablation study
In this section we perform an ablation study on (a) the tightening of intermediate bounds, (b) the new split selection heuristics, and (c) the pre-emptive shortcuts. The results can be seen in Figures 6 and 7, where we measure the approximation ratio over time for the three image datasets with both types of networks and random patches, as before. We average the results after removing any task that finishes after the first iteration (since the improvements do not matter in that case).
For the under-approximations in Figure 6 the new heuristics offer the best improvements, but the shortcuts are also very helpful. Tightening of intermediate layer bounds has a minor effect, since some of the benefit is countered by the additional computational overhead, especially for CNNs. Hence, for larger networks (where LiRPA slows down) it might be beneficial to disable this improvement.
Figure 6: Comparing the average under-approximation ratio (higher is better) over time, with different improvements enabled or disabled.
For the over-approximations in Figure 7, we observe similar trends regarding the new heuristics and shortcuts providing the best improvements. The over-approximation results also contained more instances that finish after one iteration, which indicates that computing an over-approximation is easier.
Figure 7: Comparing the average inverse over-approximation ratio (higher is better) over time, with different improvements enabled or disabled.
6 Use cases
In this section we demonstrate a range of use cases for PREMAP, showing how to exploit over- and under-approximations. In Section 6.1, we extend the discussion of the leading example of certification against patch attacks by demonstrating how to use PREMAP with a wider range of patches. In Section 6.2 we extend the outlier detection use case from [14] showing how, with PREMAP, we are able to certify more properties than before by using both under- and over-approximations. Finally, in Section 6.3 we investigate a novel use case for PREMAP in explainable artificial intelligence (XAI). XAI aims to provide analysis tools to aid in model development and debugging [40]. We use PREMAP to highlight pixels that are important for the correct classification of a given image.
6.1 Patch attack robustness
In most experiments in this paper the patches are rectangular and cover all possible pixel values. However, PREMAP is flexible enough to handle a wider range of patch shapes. Figure 8 demonstrates a patch inspired by the sunbeam in Figure 1 (middle). We use a stop sign from the GTSRB [19] dataset. Since a sunbeam only adds brightness we define the lower bound of the input domain, $\mathbb{I}$, to be the selected image and the upper bound to contain a patch with maximum possible brightness (white) of an irregular shape (pixelated triangle).
Figure 8: Certifying robustness against a non-rectangular patch. The first two images show the bounds of the input domain, $\mathbb{I}$, and the last one (right) shows progress of PREMAP over time.
Figure 8 (right) also demonstrates the anytime property of the PREMAP algorithm: if we stop early we still get an under-approximation, just a smaller one, and over time more approximations approach the $0.9$ threshold. The curve is also characteristic of PREMAP. There is an initial delay before the bounds become tight enough to certify any coverage and the improvements slow down as the subdomains reduce in size.
6.2 Out-of-distribution detection
In [14] the authors propose that preimage over-approximations can be used for certified out-of-distribution (OOD) detection. During training an additional logit is added to represent the out-of-distribution data that is added to the training data [41]. Comparing that logit with the logits for the real classes lets us detect outliers.
The over-approximations of INVPROP [14] divide the input-space into certified outliers and probable non-outliers. With PREMAP we are able to add a new category: certified non-outliers. This is due to PREMAP being able to compute both under- and over-approximations.
Figure 9: Using preimages for certifying out-of-distribution detection.
In Figure 9 we use the same model as in [14], a fully connected neural network with layer sizes $200$, $200$, and $3$. The model was trained on non-outliers from two, spatially separated, normal distributions. The outliers were generated by uniformly sampling from the input space, $2,2]^{2}$, and removing any sample too close to a non-outlier. With PREMAP we get a $0.9$ under-approximation of the non-outliers in $6.9$ seconds and a $1.1$ over-approximation in $2.9$ seconds when running only on the CPU (implemented as approximations of the outliers with thresholds $1.02$ and $0.98$). For reference, in [14] INVPROP achieved an over-approximation of $1.14$ in $12.0$ seconds utilizing a GPU.
6.3 Explainability (XAI)
A popular approach in XAI is to cover part of an image and see if the classification changes [42]. The reasoning is easy to understand: if the prediction changes that area must have been important. However, this requires careful design of the replacement pattern [43]. By using PREMAP we do not consider any singular replacement pattern, but rather all possible values in the preimage. We show both a qualitative example in Figure 10 and a more thorough evaluation in Figure 11, including a comparison in Figure 12.
Figure 10: Using PREMAP to investigate which parts of the image are important for classification. The figures are, from left to right, the original image, the estimated preimage volume (higher means more robust), the approximation ratio of the preimage, and the number of iterations (lower means easier).
In Figure 10 we apply $4\times 4$ patches on a grid covering the image on the left. In the second image from the left we show the estimated volume of the preimage (not the approximation) for every patch location. By using PREMAP we can quantify the proportion of all possible patches that would change the prediction, which lets us rank the importance of different parts of the image. Here we see a patch location near the ‘5’ being highlighted.
Furthermore, in the two images on the right of Figure 10, we plot the approximation ratio and number of iterations before the 10-minute time limit. More areas around the ‘5’ are highlighted, indicating that the model is highly non-linear in those areas [44], which makes sense as other speed limit signs only differ in the first digit.
To evaluate multiple explanations we follow [45] and create a dataset where we combine one labelled image (from GTSRB) with three unrelated images (from CIFAR-10) and train a CNN classifier. The leftmost image of Figure 11 shows an example. This lets us control the location of the meaningful part of the image: a random quadrant during training and top left for evaluation.
Figure 11: Aggregating multiple explanations on a dataset combining four images. The figures are (from left to right): An example, the average preimage volume, the average approximation ratio, and the (extrapolated) time required to reach an approximation ratio of $0.9$.
In the second image of Figure 11 we visualize the mean preimage volume from multiple explanations. All the highlighted patches are in the correct quadrant. However, in roughly one fifth of all explanations we find no patch able to change the prediction. In these cases we might want a secondary indicator of importance, such as the time required by PREMAP. When the time limit is exceeded, we use the achieved approximation ratio to extrapolate the time required to reach the threshold of $0.9$ (shown in the rightmost image).
In Figure 12 we sort the pixels according to the importance and count the fraction of pixels that are within the correct quadrant for increasing percentiles. For comparison, we also include three dedicated XAI methods. We use the partition version of SHAP [46] with $1000$ samples, which produces patches of similar size to what we use for PREMAP, and “inpaint_telea” as the replacement pattern. For LIME [47] we use manual “superpixels” that match the grid we used for PREMAP and the default grey replacement. Finally, we use LRP [48, 49] that produces pixel-level importance values.
Figure 12: Fraction of the top highlighted pixels in the right quadrant. Higher is better below 25% (the size of one quadrant).
The pixels highlighted due to the preimage volumes are always in the correct quadrant, but some explanations lack any rankings. If we use time as a secondary criterion to rank pixels with equally sized preimages, the results are close to the best XAI method, SHAP. The LRP explanations are highly granular, often focusing on edges, so LRP fares less well in this evaluation and LIME performs even worse.
The dedicated XAI methods are faster, but the preimage based rankings offer not only an importance score but also quantify how sensitive the prediction is in different parts of the image.
7 Conclusions
The contributions of this work were twofold: to develop algorithmic extensions to the PREMAP algorithm for finding preimage approximations for neural networks, and to demonstrate its usability through a range of use cases. We have carefully devised a set of heuristics for selecting splits, improved bound tightening, and implemented efficiency improvements that systematically outperform the original PREMAP on both smaller control benchmarks and patch attack certification on images. We also demonstrated how the ability to compute both under- and over-approximations can be utilized in certification of adversarial robustness and outlier detection.
Future work includes applying PREMAP to other domains where certification is needed, for example biometric security or medical diagnosis. Furthermore, extending PREMAP to non-uniform input spaces would unlock additional applications, for example, prioritizing dangerous conditions.
Appendix A Heuristics
Tables V and VI show the heuristic selection results similar to Table II but with under- and over-approximations separated (using the same hyperparameters of $\alpha=8$, $\gamma=1.0$, and a Laplacian kernel). Both tables show similar results to Table II in Section 5.2, prioritizing the same heuristics. The over-approximations show lower predicted $\delta$, which means that the choice of heuristic has a lesser impact on the results. This is unsurprising considering the evidence in Tables III and 5 showing over-approximations being easier.
TABLE V: Combinations of heuristics for under-approximations
$\boldsymbol{\delta}$ Weights
$0.811$ area: $1.00$
$0.770$ under: $1.00$
$0.748$ extra: $1.00$
$0.558$ gap: $1.00$
$0.870$ area: $1.00$, gap: $0.75$
$0.869$ extra: $1.00$, area: $1.00$
$0.864$ area: $1.00$, gap: $0.50$
$0.861$ extra: $1.00$, area: $0.75$
$0.899$ area: $1.00$, extra: $0.75$, gap: $0.50$
$0.895$ extra: $1.00$, area: $1.00$, gap: $0.50$
$0.889$ extra: $1.00$, area: $1.00$, gap: $0.25$
$0.887$ area: $1.00$, extra: $0.50$, gap: $0.50$
$0.893$ area: $1.00$, extra: $0.75$, gap: $0.50$, under: $0.25$
$0.892$ area: $1.00$, under: $0.75$, gap: $0.50$, soft: $0.50$
$0.891$ area: $1.00$, under: $0.75$, gap: $0.75$, soft: $0.50$
$0.891$ area: $1.00$, under: $0.75$, soft: $0.75$, gap: $0.50$
$0.889$ area: $1.00$, under: $0.75$, gap: $0.50$, soft: $0.50$, width: $0.25$
$0.888$ under: $1.00$, area: $0.75$, gap: $0.50$, soft: $0.50$, extra: $0.25$
$0.888$ area: $1.00$, under: $0.75$, gap: $0.50$, soft: $0.50$, extra: $0.25$
$0.887$ under: $1.00$, area: $0.75$, soft: $0.75$, gap: $0.50$, extra: $0.25$
TABLE VI: Combinations of heuristics for over-approximations
$\boldsymbol{\delta}$ Weights
$0.608$ under: $1.00$
$0.585$ extra: $1.00$
$0.523$ soft: $1.00$
$0.514$ loose: $1.00$
$0.675$ extra: $1.00$, area: $0.75$
$0.652$ extra: $1.00$, area: $1.00$
$0.649$ extra: $1.00$, under: $1.00$
$0.645$ extra: $1.00$, area: $0.50$
$0.696$ extra: $1.00$, area: $0.75$, under: $0.50$
$0.693$ extra: $1.00$, under: $0.75$, area: $0.75$
$0.690$ under: $1.00$, area: $0.75$, loose: $0.75$
$0.689$ extra: $1.00$, area: $0.75$, soft: $0.25$
$0.709$ extra: $1.00$, under: $0.75$, area: $0.75$, gap: $0.25$
$0.706$ extra: $1.00$, area: $0.75$, under: $0.50$, gap: $0.25$
$0.702$ extra: $1.00$, under: $0.75$, area: $0.50$, gap: $0.25$
$0.700$ extra: $1.00$, area: $0.75$, under: $0.50$, soft: $0.25$
$0.700$ extra: $1.00$, under: $0.75$, area: $0.75$, gap: $0.25$, balance: $0.25$
$0.699$ extra: $1.00$, under: $0.75$, area: $0.75$, gap: $0.25$, loose: $0.25$
$0.698$ extra: $1.00$, area: $0.75$, under: $0.50$, gap: $0.25$, soft: $0.25$
$0.698$ extra: $1.00$, area: $0.75$, under: $0.50$, gap: $0.25$, balance: $0.25$
Appendix B Neural networks
For the image datasets we use two types of neural networks: fully connected and convolutional. The architectures are described in Table VII. The networks were trained with cross-entropy loss using the AdamW [50] optimizer with brightness and hue variation until the accuracy on the test set stopped decreasing (usually around 50 epochs). For the reinforcement learning tasks we used the same networks as in [16] and for the out-of-distribution detection we used the same network as in [14].
TABLE VII: Architectures for the neural networks
Fully connected Convolutional
Layer Output size Layer Output size
Input $32\times 32\times 3$ Input $32\times 32\times 3$
Flatten $3072$ Convolutional $30\times 30\times 3$
Linear $300$ ReLU $30\times 30\times 32$
ReLU $300$ Average pooling $15\times 15\times 32$
Linear $300$ Convolutional $12\times 12\times 32$
ReLU $300$ ReLU $12\times 12\times 32$
Linear $300$ Average pooling $6\times 6\times 32$
ReLU $300$ Convolutional $4\times 4\times 64$
Dropout $300$ Average pooling $2\times 2\times 64$
Linear $|\mathbf{y}|$ ReLU $2\times 2\times 64$
Flatten $256$
Dropout $256$
Linear $|\mathbf{y}|$
Acknowledgments
This project received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (FUN2MODEL, grant agreement No. 834115) and ELSA: European Lighthouse on Secure and Safe AI project (grant agreement No. 101070617 under UK guarantee). MZ worked on this project during an internship at the University of Oxford.